# Counterfactual Simulatability of LLM Explanations for Generation Tasks

### v1

@@ 4.4 Experimental setup
Figure 1 provides a high-level overview of our evaluation framework. First, using randomly sampled inputs from our dataset, we prompt GPT-4 Turbo (see section 4.2), to generate outputs and explanations. For each of these explanations, we use GPT-4 Turbo to generate 3 relevant counterfactuals, providing the explanation in the prompt as context. Then, we use GPT-4 Turbo to decompose the explanation into atomic units (i.e., $e_{x}\rightarrow\{a\}$) and, in the medical suggestion use case, instruct it to extract units into two groups for patient details or suggested actions. We opted for GPT-4 Turbo as the LLM in our evaluation framework due to its performance and relatively low cost. Finally, we generate counterfactual outputs using the same prompt as the first step and have humans examine the counterfactual and counterfactual output to annotate simulatability and precision for each atomic unit of the explanation. To evaluate the ability of GPT-4 Turbo to break down explanations, we also instruct the annotators to indicate whether each unit was extracted correctly and whether any details that should have been extracted are missing. Furthermore, we investigate if increasing the number of counterfactuals generated per explanation affects generality, but find no noticeable differences (appendix E). Additional details for the human evaluation and the prompts used are provided in appendix A and F respectively.
Comparison of human-LLM annotations.
We evaluate whether using an LLM to automate human annotation is feasible by prompting GPT-4 Turbo with similar instructions and assessing the inter-annotator agreements between human-human and human-LLM pairs.
Automatic evaluation.
Finally, using GPT-4 Turbo for annotation, we repeat our framework and conduct a larger evaluation of three popular LLMs using more data and five counterfactuals per explanation. Additional details for the automatic evaluation are provided in appendix B.
5 Results
5.1 Intermediate results
Annotators find that the LLM successfully breaks down explanations with 96% and 57% accuracy for news summarization and medical suggestion respectively. For medical suggestion errors, the LLM either fails to identify a key detail or extracts an atomic unit incorrectly (extracted a unit it should not have or classified it as the incorrect type of information). A detailed break down of parsing errors is provided in appendix C.
While almost all ($74/76$) generated counterfactuals are deemed simulatable in the summarization setting, only slightly more than half are for medical suggestion ($52/90$). Most non-simulatable counterfactuals for medical suggestion contain $0.4-0.8$ of the atomic units of the explanation, indicating that counterfactual generation is more difficult in this setting. The distribution of generated counterfactuals, bucketed by the proportion of atomic units that appear is provided in appendix D.
5.2 Human evaluation
Task Explanation Method Generality Precision
News Summarization Chain of Thought 0.52 0.81
Post Hoc 0.49 0.89
Medical Suggestion Chain of Thought 0.20 0.51
Post Hoc 0.26 0.59
Table 1: Generality and precision across explanation types and tasks for human evaluation.
GPT-4 Turbo explanations lead to generalizable counterfactuals and consistent mental models in the case of summarization, with approximately $0.8$ of the inferred information appearing in the counterfactual output on average. In contrast, explanations in the medical suggestion setting are less generalizable and less precise (approximately $0.5$), indicating that LLMs may struggle more to reliably explain their behavior for this task. Chain-of-thought and Post-hoc explanations lead to similar results (see table 1).
Additionally, as a sanity check, we generate counterfactual outputs conditioned on the original explanation and verify that these result in a precision score of $1.00$. Since we expect that the model will perfectly follow the explanation, this precision score indicates that our framework does indeed capture how well models follow their explanations. Thus, the low precision scores for medical suggestion are due to the model’s behavior not adhering to its explanation on that task rather than the evaluation setup.
5.3 Automatic evaluation
@@
GPT-4 Turbo is able to approximate human annotation for our tasks.
Table 2 reports the pairwise Cohen’s Kappa between each pair of annotators and the LLM. We find that overall, GPT-4 Turbo achieves similar inter-annotator agreement to humans, with an average of 0.61 for human-LLM pairs compared to 0.54 for human-human pairs in news summarization and 0.65 for human-LLM pairs compared to 0.74 for human-human pairs in medical suggestion.
@@
News Summarization Claude 3.7 Sonnet 48 189 0.67 0.93 45 173 0.62 0.84
GPT-4 46 160 0.59 0.84 48 193 0.65 0.78
Llama 3 47 172 0.67 0.74 43 153 0.67 0.66
Medical Suggestion Claude 3.7 Sonnet 24 55 0.21 0.48 24 78 0.20 0.66
GPT-4 36 103 0.20 0.46 36 122 0.19 0.65
Llama 3 24 69 0.19 0.56 30 110 0.20 0.66
Table 3: Generality and precision across models, explanation types, and tasks for automatic evaluation.
The results of the automatic evaluation, presented in table 3, further support our findings from the human evaluation. Namely, models are much better able to accurately explain their behavior for summarization as opposed to medical suggestion while remaining general such that these explanations apply to a diverse set of counterfactuals. Chain-of-thought explanations tend to lead to better mental models for summarization, which may reflect the skill-based nature of this task. On the other hand, Post-hoc explanations lead to more precise explanations for medical suggestion. Models may also differ in their ability to describe their behavior in a generative setting, for instance, Claude 3.7 Sonnet demonstrates noticeably better precision scores for news summarization compared to other models.
@@ 6 Discussion
LLM explanations are helpful for skill-based generation tasks but struggle for knowledge-based generation tasks.
While model explanations do enable users to infer pieces of information that will appear in counterfactual outputs in both settings, the mental models they produce are more accurate in summarization compared to medical suggestion. Furthermore, summarization explanations, employing a high-level approach in their explanation, also lead to counterfactuals that are more generalizable. These may suggest that LLMs are better able to describe their behavior for skill-based tasks, where users can reliably infer the elements that will appear in the outputs, compared to knowledge-based tasks, where users are less capable of reliably inferring specific points (e.g., suggestions). Additionally, using Chain-of-thought prompting, which aligns with the skill-based nature of summarization, may lead to more precise explanations.
These findings may reflect the predictability of model behavior on different tasks types. For medical question, minor alterations to the question may lead to very different answers. For instance, if a user describes ”I experience chest pain sometimes when I exercise.” a model might respond ”consider reducing exercise intensity, try warming up thoroughly before exercise …”. However, if a user changes their question slightly to ”I experience chest pain sometimes when I exercise. I started taking a new pre-workout supplement with high caffeine recently.”, although they are expressing the same symptoms, because of the presence of additional information the model might respond ”discontinue the supplement or use smaller dosages, the chest pain is likely related to caffeine-induced heart palpitations …”. These differences in the sensitivity of the model’s output relative to the input may lead to the differences in precision and generality observed in our experiments.
@@
Although LLMs are able to automate the human annotation steps in our evaluation, they demonstrate issues in other aspects for the medical suggestion task. Specifically, annotators identify many errors in the parsing of explanations into atomic units, where the LLM misses key information or mis-classifies the information. In one instance, “possible heat-induced asthma” is incorrectly extracted as an expressed symptom when in reality it is a potential cause. Unlike the summarization setting, this additional requirement of classifying the type of information introduces more complexity, making the LLM a less effective tool in the evaluation pipeline. Additionally, we found that the LLM is unable to produce as many simulatable counterfactuals in the medical domain compared to summarization. There is significant room for improvement towards adapting counterfactual simulatability for knowledge-based tasks like medical suggestion, where the explanation and decision process for an LLM is less explicit and relies heavily on knowledge encoded in the model.
Appendix A Human Evaluation
Annotators: Two groups of 3 student annotators each from the Columbia NLP lab were recruited for the human evaluation (one group per task). Since these tasks were meant to replicate LLMs in a practical setting, where a general user is asking for a brief summary of an article or seeking some quick medical advice, the annotators were not required to have any specialized domain knowledge.
Distribution of Annotations: For each task and explanation type (Chain-of-thought and Post-hoc), annotators were asked to annotate across a total of 15 explanations with 3 counterfactuals each, resulting in 45 samples (explanation, counterfactual pairs). In order to compare human-human agreement against human-LLM agreement, each annotator was assigned the same set of 3 samples plus an additional 4. Not all annotations were completed for news summarization, leading to a slight discrepancy in annotations across tasks.
News Summarization Instructions:
  • Parsed Explanation Annotations: Check if the explanation (Column B) is parsed correctly in Column C. If yes, put ”ok”. If not, write which information in the explanation is missing / which info in column C are hallucinating.
  • Document Annotations: Check whether each item in the parsed explanation appears in the document or not. Put either ”no” or ”yes”.
  • Summary Annotations: Check whether each item in the parsed explanation appears in the summary or not. Put either ”no” or ”yes”
Medical Suggestion Instructions:
Tasks 1, 3:
  • for each of the ”Extracted Points”, label ”Y” or ”N” if the point was parsed correctly from the AI explanation
Tasks 2, 4:
  • for each of the ”Extracted Points”, label ”Y” or ”N” if the point is reflected in the counterfactual (task 2) or AI suggestion (task 4)
  • the matching is flexible, i.e., point does not necessarily need to appear in exactly the same wording to be marked ”Y”
  • for example, in example 3 the point mentions ””hemoglobin level of 10.5, slightly below normal”” these specific scores do not need to be mentioned, rather ””below normal levels”” is sufficient.
Appendix B Automatic Evaluation
The automatic evaluation consisted of 50 explanations per task and generating 5 counterfactuals each (250 explanation, counterfactual pairs total). The models evaluated were: Anthropic’s Claude 3.7 Sonnet, OpenAI’s GPT-4, and Meta’s Llama 3.3 70B Instruct. These models were chosen as they represent a few popular LLM options as well as a mix of proprietary and open-source models.
Appendix C Explanation Parsing Quality Results
We show a categorized break down of the reported errors for explanation parsing in both tasks.
@@ Appendix C Explanation Parsing Quality Results
Table 4: Observed errors in explanation parsing across tasks.
Appendix D Counterfactual Generation: Simulatability
We calculate the proportion of atomic units that appear in each counterfactual for the human evaluation (where 1.00 indicates a simulatable counterfactual), bucket them, and display the results below.
@@ Appendix D Counterfactual Generation: Simulatability
Appendix E Counterfactual Generation: Generality
We show how generality and the number of simulatable counterfactuals changes as we adjust the number of counterfactuals generated per explanation. For news summarization and 10 counterfactuals per explanation, not all 100 counterfactuals were generated due to errors parsing the LLMs output. Since news summarizations are more expensive to produce (longer context prompt), we only use 10 explanations compared to 30 for medical suggestion.
@@ Appendix E Counterfactual Generation: Generality
Table 6: Generality and simulatability as the number of generated counterfactuals per explanation increases.
Appendix F LLM Prompts

### v2

@@ 4.5 Experimental setup
We run two main experiments: (1) a human evaluation where GPT-4 Turbo explanations are evaluated using human annotations and (2) a larger-scale, fully automatic evaluation of multiple models (Claude 3.7 Sonnet, GPT-4, Llama 3) using GPT-4 Turbo as the annotator. This allows us to assess the feasibility of using an LLM as an annotator before adopting the automatic approach that can scale to larger experiments.
@@
Using randomly sampled inputs from our dataset, we prompt GPT-4 Turbo to generate outputs and explanations. Note that here, GPT-4 Turbo is both the model being evaluated and the model being used in the pipeline. For each explanation, we use GPT-4 Turbo to generate 3 relevant counterfactuals, providing the explanation in the context. Then, we use GPT-4 Turbo to decompose the explanation into atomic units (i.e., $e_{x}\rightarrow\{a\}$) and, in the medical suggestion use case, instruct it to extract units into two groups for patient details or suggested actions. Finally, we generate counterfactual outputs using the same prompt as the first step and have humans examine the counterfactual and counterfactual output to annotate simulatability and precision for each atomic unit of the explanation.
Additionally, we assess the ability of GPT-4 Turbo to break down explanations by instructing the annotators to indicate whether each unit was extracted correctly and whether details that should have been extracted are missing. We also briefly investigate if increasing the number of counterfactuals generated per explanation affects generality, but find no noticeable differences (Appendix C).
@@
We assigned 3 student annotators per task to evaluate 15 explanations along with 3 generated counterfactuals each (45 explanation, counterfactual pairs total). Each annotator was given an overlapping set of 3 explanations (to measure human-human agreement) plus an additional 4. Since they annotated the atomic units of each explanation, the exact number of annotations varied but approximately resulted in 260 annotations per person (see Table 2). The annotators were not required to have any specialized domain knowledge. One annotator (news summarization annotator #3 in Table 1) did not complete all annotations, leading to a slight discrepancy in annotations across tasks. The prompts and annotations instructions are provided in Appendix D and E respectively.
Comparison of human-LLM annotations.
We evaluate the feasibility of using an LLM to automate human annotation by prompting GPT-4 Turbo with similar instructions and measuring the inter-annotator agreements between human-human and human-LLM pairs.
Automatic evaluation.
Finally, using GPT-4 Turbo as the annotator, we scale up our experiment to evaluate three LLMs using more data and generating five counterfactuals per explanation. For the automatic evaluation, 50 explanations along with 5 generated counterfactuals each (250 explanation, counterfactual pairs total) were evaluated.
5 Results
5.1 Intermediate results
Annotators find that the LLM successfully breaks down explanations with 96% and 57% accuracy for news summarization and medical suggestion respectively. For medical suggestion errors, the LLM either fails to identify a key detail or extracts an atomic unit incorrectly (extracted a unit it should not have or classified it as the incorrect type of information). Further investigation reveals that errors often involve minor details whereas the main points are successfully extracted (see Appendix A).
While almost all ($74/76$) generated counterfactuals are deemed simulatable in the summarization setting, only slightly more than half are for medical suggestion ($52/90$). Most of these non-simulatable counterfactuals contain 0.40.80.4-0.8 of the atomic units of the explanation, indicating that counterfactual generation is more challenging in this setting. The distribution of generated counterfactuals, sorted by the proportion of atomic units that appear is shown in Appendix B.
5.2 Human evaluation
GPT-4 Turbo explanations lead to generalizable counterfactuals and consistent mental models in the case of summarization, with approximately $0.8$ of the inferred information appearing in the counterfactual output. In contrast, explanations in the medical suggestion setting are less generalizable and less precise (approximately $0.5$), indicating that LLMs may struggle more to reliably explain their behavior for this task. Chain-of-thought and Post-hoc explanations lead to similar results (see Table 2). For each pair of settings, we fit an independent sample t-test and find that the difference in both metrics when compared across tasks are significant ($p<0.05$), but not when compared across explanation types within the same task.
Task Explanation Generality Precision
News
Summarization
Chain-of-
thought
0.52 0.81
Post-hoc 0.49 0.89
Medical
Suggestion
Chain-of-
thought
0.20 0.51
Post-hoc 0.26 0.59
Table 2: Generality and precision scores for GPT-4 Turbo from the human evaluation.
As a sanity check for the pipeline, we generate counterfactual outputs conditioned on the original explanation and verify that these result in a precision score of $1.00$. This validates that our framework captures how well models follow their explanations. Therefore, the low precision scores for medical suggestion are due to the model’s behavior not adhering to its explanation on that task rather than the evaluation setup.
@@ 5.2 Human evaluation
News
Summarization
Claude 3.7 Sonnet
48 189 0.67 0.93 45 173 0.62 0.84
GPT-4
46 160 0.59 0.84 48 193 0.65 0.78
Llama 3
47 172 0.67 0.74 43 153 0.67 0.66
Medical
Suggestion
Claude 3.7 Sonnet
24 55 0.21 0.48 24 78 0.20 0.66
GPT-4
36 103 0.20 0.46 36 122 0.19 0.65
Llama 3
24 69 0.19 0.56 30 110 0.20 0.66
Table 3: Generality and precision metrics across models, explanation types, and tasks from the automatic evaluation.
5.3 Automatic evaluation
GPT-4 Turbo is able to approximate human annotation for our tasks.
Table 1 reports the pairwise Cohen’s Kappa between each pair of annotators and the LLM. We find that overall, GPT-4 Turbo achieves similar inter-annotator agreement to humans, with an average of 0.61 for human-LLM pairs compared to 0.54 for human-human pairs in news summarization and 0.65 for human-LLM pairs compared to 0.74 for human-human pairs in medical suggestion. We calculate a two-sided p-value for ratings between all annotation pairs and find that the observed agreement is significant.
The results of the automatic evaluation, presented in Table 3, are in line with our findings from the human evaluation. Namely, models are better at accurately explaining their behavior for summarization compared to medical suggestion while also remaining general such that these explanations apply to diverse counterfactuals. Chain-of-thought explanations lead to better mental models for summarization, which may reflect the skill-based nature of this task. On the other hand, Post-hoc explanations lead to more precise explanations for medical suggestion. Models may also differ in their ability to describe their behavior in a generative setting, for instance, Claude 3.7 Sonnet demonstrates noticeably better precision scores for news summarization compared to other models.
@@ 6 Discussion
LLM explanations can be helpful for skill-based generation tasks but may struggle for knowledge-based generation tasks.
While model explanations do enable users to infer pieces of information that will appear in counterfactual outputs in both settings, the mental models produced are more accurate for summarization compared to medical suggestion. Furthermore, summarization explanations, which employ a high-level approach to their explanation, are also more generalizable. This may suggest that LLMs are better able to describe their behavior for skill-based tasks, where users can reliably infer the elements that will appear in the outputs, compared to knowledge-based tasks, where users are less capable of inferring specific points (e.g., suggestions). Additionally, utilizing Chain-of-thought prompting, which aligns with the skill-based nature of summarization, may lead to more precise explanations. It is important to consider that our evaluation only examined news summarization and medical suggestion specifically. As such, the results may vary for other skill-based and knowledge-based tasks.
These findings may reflect the predictability of model behavior on different tasks types. For medical suggestion, minor variances in the question may lead to very different answers. For instance, if a user describes “I experience chest pain sometimes when I exercise.” a model might respond “consider reducing exercise intensity, try warming up thoroughly before exercise …”. However, if a user changes their question slightly to “I experience chest pain sometimes when I exercise. I started taking a new pre-workout supplement with high caffeine recently.”, although they are expressing the same symptoms, because of the presence of additional information the model might respond “discontinue the supplement or use smaller dosages, the chest pain is likely related to caffeine-induced heart palpitations …”. These differences in the sensitivity of the model’s output relative to the input may lead to the observed differences in precision and generality in our experiments.
@@
Although LLMs are able to automate the human annotation steps in our evaluation, they demonstrate issues in other aspects for the medical suggestion task. Specifically, our annotators identify many errors in the parsing of explanations into atomic units, where the LLM misses key information or mis-classifies the information. In one instance, “possible heat-induced asthma” is incorrectly extracted as a key symptom when in reality it is a potential cause. Unlike the summarization setting, this added requirement of classifying the type of information introduces more complexity, making the LLM a less effective tool in the evaluation pipeline. Additionally, we found that the LLM is unable to produce as many simulatable counterfactuals in the medical domain compared to summarization. There is significant room for improvement towards adapting counterfactual simulatability for knowledge-based tasks like medical suggestion, where the explanation and decision process for an LLM is less explicit and relies heavily on knowledge encoded in the model.
7 Limitations
First, the generality metric relies on our LLM (GPT-4 Turbo) being able to generate diverse, yet simulatable counterfactuals. Although we briefly experimented with counterfactual generation, more work can be done to assess whether other models or prompting strategies may lead to higher quality counterfactuals. Additionally, cosine similarity might not be the optimal way to judge diversity across different tasks as it is heavily content-based. For example, while many counterfactuals for summarization followed the same structure (e.g., a headline followed by when/where it occurred followed by a quote …), they were still rewarded with high generality because the specific topics and words differed.
Second, in a few cases there was ambiguity when determining whether atomic units were ‘present’in the counterfactual/the counterfactual output. Since these units of information may appear with different wording or even be implied, we encouraged annotators to use their best judgement and leave notes. We leave investigation into ways of improving this matching to future work.
Finally, we focused on demonstrating our framework using explanations generated from a few popular LLMs. Testing newer types of models (e.g., reasoning models) and different task setups (e.g., multiple interactions) are a promising direction. Furthermore, expanding the set of metrics to consider aspects such as explanation comprehensiveness, faithfulness, and plausibility, and assessing how they vary with counterfactual simulatability, may also be worthwhile.
Supplementary Materials Availability Statement:
Our source code as well as an anonymized version of the data used in the human evaluation is made available at github.com/mlimpijankit/counterfactual-simulatability-generation-tasks.
8 Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005 and the National Science Foundation and by DoD OUSD (R&E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence). The views, opinions and/or findings expressed are those of the author and should not be interpreted as representing the official views or policies of the Department of Defense, the National Science Foundation or the U.S. Government. One of the authors holds an equity interest in OpenAI.
Appendix A Explanation Decomposition
We categorize the errors reported by our annotators for explanation decomposition using GPT-4 Turbo during the human evaluation. Examples of each type of error are also provided below. Note that errors are subjective and often involve small details.
@@ Appendix A Explanation Decomposition
Table 4: Explanation parsing errors across tasks.
Example of a Missing Extraction
In the example below, “anxiety and stress” are mentioned in the explanation but not extracted.
Explanation.
The symptoms you describe, including chest tightness, difficulty breathing, and the presence of white phlegm, along with discomfort in the stomach and worsening symptoms throughout the day, suggest a complex interaction between gastrointestinal and possibly psychosomatic factors. Superficial gastritis diagnosed previously might not fully explain the severity and range of your symptoms, especially if the treatment provided little relief. The persistence and nature of your symptoms indicate the need for a more comprehensive gastrointestinal evaluation to rule out other conditions such as gastroesophageal reflux disease (GERD), which can cause acid reflux, chest discomfort, and respiratory symptoms. Additionally, the significant impact of anxiety and stress on your physical health suggests that addressing these psychological factors is crucial. Cognitive behavioral therapy (CBT) is effective in managing anxiety and may reduce the intensity of the symptoms by helping you develop better coping mechanisms.
GPT-4 Turbo Parsed Explanation (Symptoms).
  • chest tightness and difficulty breathing
  • presence of white phlegm
  • discomfort in the stomach
  • symptoms worsen throughout the day
  • previously diagnosed with superficial gastritis
  • treatment for gastritis provided little relief
  • Missing Extraction: anxiety and stress
GPT-4 Turbo Parsed Explanation (Suggestions).
  • conduct a comprehensive gastrointestinal evaluation to rule out conditions like GERD
  • address psychological factors such as anxiety and stress
  • consider Cognitive Behavioral Therapy (CBT) to manage anxiety and develop coping mechanisms
Example of an Incorrect Extraction
In the example below, “possible heat-induced asthma or heat intolerance” is incorrectly extracted in the symptoms category. The annotator noted that this was more of a potential cause rather than an expressed symptom.
Explanation.
The patient experiences difficulty breathing, particularly during the summer or in hot temperatures, which improves significantly in cooler environments such as air-conditioned rooms or when using a fan. This pattern suggests that the symptoms may be related to heat-induced asthma or a condition known as heat intolerance. Heat can lead to increased airway resistance and trigger respiratory symptoms in sensitive individuals. It is also possible that environmental allergens, which are often more prevalent in warm weather, could be exacerbating the condition. To manage these symptoms, staying cool and avoiding heat exposure is crucial. However, a thorough evaluation by a healthcare professional is necessary to rule out other respiratory conditions and to confirm the diagnosis.
GPT-4 Turbo Parsed Explanation (Symptoms).
  • difficulty breathing, especially during summer or in hot temperatures
  • improvement in symptoms in cooler environments
  • possible heat-induced asthma or heat intolerance (Incorrect Extraction)
GPT-4 Turbo Parsed Explanation (Suggestions).
  • stay cool and avoid heat exposure
  • undergo a thorough evaluation by a healthcare professional to rule out other respiratory conditions and confirm the diagnosis
Appendix B Counterfactual Generation Simulatability
For each counterfactual in the human evaluation, we measure the extent to which they are simulatable by the proportion of atomic explanation units they contain. This distribution is displayed in Table 5.
@@ Appendix B Counterfactual Generation Simulatability
Appendix C Counterfactual Generation Generality
We assess how generality and simulatability changes as the number of counterfactuals generated per explanation is increased. Since news summarizations are expensive to generate due to the prompt length, we only use 10 explanations in this experiment compared to 30 for medical suggestion. Note that in the summarization and 10 counterfactuals setting, not all 100 counterfactuals were generated due to errors parsing the LLMs output.
@@ Appendix C Counterfactual Generation Generality
Table 6: Generality and simulatability metrics as the number of generated counterfactuals per explanation increases.
Appendix D LLM Prompts